OpenStudy (anonymous):

I don't understand. Each year, Terrell buys a plane ticket to visit his grandmother for a week. Last year, he paid one hundred thirty-five dollars. This year, he paid one hundred eighty-nine dollars. What was the percent increase in the cost of his plane ticket?

OpenStudy (aivantettet26):

given: $135 = first ticket The next year its $189 = second ticket

OpenStudy (aivantettet26):

Solution:

OpenStudy (aivantettet26):

135 is 100 percent of the ticket

OpenStudy (aivantettet26):

189 - 135 = 54

OpenStudy (aivantettet26):

54 is what percent of 135/?

OpenStudy (aivantettet26):

you there?

OpenStudy (anonymous):

45?

OpenStudy (aivantettet26):

40 :)

OpenStudy (anonymous):

o thank you.:)

OpenStudy (aivantettet26):

54/135 x 100 = 40

OpenStudy (aivantettet26):

to check if it is correct 135 - 54 = 81 81/135 x 100 = 60

OpenStudy (aivantettet26):

60 + 40 = 100 percent :)

OpenStudy (anonymous):

ok i will try that on my next one